5 amp . current is passing through a dry cell for 2 hours .the value produced electric current will be ?give explanation

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
dear student,
q=it=2×3600×5=7200×5=36000C

q=3.6×10^4
thank you